A ball is thrown upward from the roof of a building 100 m tall with an initial velocity of 20 m/s.
When will the ball reach a height of 80m?
I would appreciate help with this one. Thank you!

Answer by Edwin McCravy(20086)

A ball is thrown upward from the roof of a building 100 m tall with an initial velocity of 20 m/s.
When will the ball reach a height of 80m?
I would appreciate help with this one. Thank you!

The formula for the height h is

                h = h0 + v0t - 4.9t²

h = the final height = 80 m
where h0 = the initial height = 100 m
v0 = the initial velocity = 20 m/s
t = the time to reach height h, so t is unknown.

                h = h0 + v0t - 4.9t²
               80 = 100 + 20t - 4.9t²
 4.9t² - 20t - 20 = 0

Solve by the quadratic formula:

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        t = —————————————
                2a 

where a = 4.9; b = -20; c = -20

                       ___________________
             -(-20) ± Ö(-20)²-4(4.9)(-20)
        t = ———————————————————————————————
                       2(4.9) 
                   _______ 
             20 ± Ö400+392
        t = ———————————————
                  9.8

                   ___ 
             20 ± Ö792
        t = ————————————
                 9.8 

                   _____ 
             20 ± Ö36·22
        t = —————————————
                9.8  

                    __ 
             20 ± 6Ö22
        t = ———————————
                9.8 

         
                       
Using the +, 

                    __ 
             20 + 6Ö22
        t = ———————————
                9.8

        t = 4.912499445 seconds
                   
Using the -, 

                    __ 
             20 - 6Ö22
        x = ———————————
                9.8

        x = -.8308667917 seconds
                       
We discard that negative answer, so the final answer is:

          4.912499445 seconds

or you would probably want to round that to 4.9 seconds.

Edwin